OpenStudy (anonymous):

Two angles are complementary if their sum is 90 degrees. if the measure of the first angle is x degrees and the measure of the second angle is (3x-2) degrees, find the measure of each angle. with steps please.

OpenStudy (mathstudent55):

Since you know that the sum of the measures is 90, and one angle is x and the other is 3x - 2, add up the measures and set the sum equal to 90. Try that and get back.

OpenStudy (anonymous):

i'm not really sure how to do that :/

OpenStudy (mathstudent55):

One angle is x the other 3x -2. the sum (meaning addition) of the angles is (x) + (3x - 2)

OpenStudy (anonymous):

x=23

OpenStudy (anonymous):

x+3x-2=90 +2 +2 =x+3x=92 = 4x=92 = /4 /4 92/4 is 23 so x = 23

OpenStudy (anonymous):

so the first angle is 23 degrees and the second angle is 67 degrees!

OpenStudy (anonymous):

you get? (:

OpenStudy (anonymous):

ah, perfect! yes i get it awesome thanks :)

OpenStudy (anonymous):

sure, hey i can do that stuff haha. what math you in?

OpenStudy (anonymous):

haha yea it's simple but for me it's so complex i'm in mat022c developmental math it's kinda like basic math

OpenStudy (anonymous):

oh hah thats good, and yeah math is rediculous, im taking trigonometry and pre calculus.

OpenStudy (mathstudent55):

Now that you know that the sum of the angles is x + 3x - 2, and you know the sum of 2 complementary angfes is 90, write the equation: x + 3x - 2 = 90

OpenStudy (mathstudent55):

Then you solve the equation for x. Once you know x, x is the measure of one angle, and 3x - 2 is the measure of the other one.

OpenStudy (anonymous):

geez that must be so hard @cory_ryan

OpenStudy (anonymous):

thank you @mathstudent55

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

lol it is for me, im struggling, its senior math. i hate it, hopefully next semester i'll take math of finance.

OpenStudy (anonymous):

which is suppose to be easier.

OpenStudy (anonymous):

hopefully things will start getting better with that. what's your major?

OpenStudy (anonymous):

lol im not really majoring in anything. their college prep classes, but im going into the navy after i graduate.(;

OpenStudy (anonymous):

because im not to sure what i want to do, and it will give me time to think about it.

OpenStudy (anonymous):

oh ok well good for you for doing what you have to do :)

OpenStudy (anonymous):

yeah...you a freshman?

OpenStudy (anonymous):

in college yes

OpenStudy (anonymous):

lol oh shish haha. you got me. how college life treating yah?(:

OpenStudy (anonymous):

sweet man it's awesome

OpenStudy (anonymous):

hah thats good, do you go to a big college?

OpenStudy (anonymous):

hey message me im bout to pass tf out haha.

OpenStudy (anonymous):

alrighty